If a recipe calls for 2 cups of soup mix to 3 cups of water and I want to use 3 cups of soup mix, what formula would I use to find out how much water to use with 3 cups of soup mix? In other words, if the ratio is 2 cups mix to 3 cups water, then 3 cups mix to _____ cups of water?
Thanks! This stuff didn't stick in my head in school.

Answer by edjones(8007)

x/3=3/2
x=9/2 cups water
.
Ed